Displaying report 1-1 of 1.
Reports until 16:06, Wednesday 20 November 2024
GQuEST General
Torrey Cullen - posted 16:06, Wednesday 20 November 2024 (12006)
Continued realignment of B111B Sleds

Continuing realignment of sleds in B111B.

-Powered SHG for the first time. Note the SHG is able to be controlled via USB on the computer, and I have a USB hub near enough for this to reach. It is not plugged in currently while I decide on a final spot for the hub. Turned on and brought to temperature via the buttons on the temperature controller. Another note that we may want to double check the optimal temperature down to the 0.001 of a degree. 

-Frequency double light is coming out of the SHG. Touched up the input and output polarizations for the SHG.

-Want to realign each AOM path + fiber. Starting with AOM fiber path 4 (as labeled on the optical fiber). Powered the AOM with the patch panel + homemade chassis successfully. Replaced the mirror after the quarter waveplate to a curved f = 50 mm, ROC = 100mm mirror. The once shifted light is sent back through the aom and then fiddle with the alignment until the twice shifted beam pops up. Efficiency is 208 uW input to the AOM, 58 uW of twice shifted headed to the fiber. This is on par with previous alignments.

-Realigned into the fiber. Output is 45 uW (78% efficiency). OFC 1 has both wavelengths of light now and is ready to be realigned.

Displaying report 1-1 of 1.